The exact amount of diluted nitric acid on a mass mof copper. The nitrogen monoxide obtained occupies a volume of 0.56l at 25°C and a pressure of 98×10^3Pa. What is the mass of copper used? What is the mass of copper nitrate obtained by evaporation of the solution?


1
Expert's answer
2021-03-01T04:41:57-0500

Use ideal gas equation and get the answer.

98×10^-6pa/9.8×10^-3= 10^4

PV= mass×R×T/molar mass


10^4×0.56=mass×0.0821×298/63.5

Mass= 97.96gm